OpenStudy (ashking1):

Find the vertex of the graph of the function. f(x) = 2x2 - 8x + 9

OpenStudy (kpop4life123):

shouldnt they tell u wats x ?

Vocaloid (vocaloid):

nope, this is a function so x can be anything start by writing parentheses around the first two terms

Vocaloid (vocaloid):

f(x) = (2x2 - 8x) + 9 then factor what's inside the parentheses

OpenStudy (danjs):

it is found right in the middle , on the intersection with the line of symmetry

OpenStudy (tkhunny):

Completing the Square - f(x) = 2x^2 - 8x + 9 = 2(x^2 - 4x + ____) + 9 - 2(____) Fill in the two blanks with the magic number. What is the magic number? Think on this (a+b)^2 = a^2 + 2ab + b^2 We have a^2. We have 2ab. How do we find b^2?

OpenStudy (danjs):

, that x value

OpenStudy (ashking1):

im sorry but im still lost

OpenStudy (ashking1):

where do i start

OpenStudy (danjs):

the x value of the vertex point, is -b/(2*a) , the y value is the function f(x) at that x value

OpenStudy (danjs):

do you need to see how to arrive there, or you can just remember it

OpenStudy (danjs):

standard quadratic y = a*x^2 + b*x + c there is also a vertex form that includes the values of the coordinates of the vertex to read off

OpenStudy (ashking1):

so -b/(2*a) is -8/4=-2

OpenStudy (danjs):

notice the standard form has all addition between terms,so some of the constant values are negative f(x) = 2x2 - 8x + 9 = 2x^2 + (-8)x + 9

OpenStudy (danjs):

a = 2 b = -8 c = 9 -b/(2a) is then +8/4

OpenStudy (ashking1):

oh so its 2

OpenStudy (danjs):

yes, then find what y is there to complete the point

OpenStudy (danjs):

x=2 is the line of symmetry for the graph

OpenStudy (danjs):

vertex at (2 , f(2) )

OpenStudy (ashking1):

give me a second sorry

OpenStudy (danjs):

here is a short sorta summary if you want to read it http://hotmath.com/hotmath_help/topics/vertex-of-a-parabola.html

OpenStudy (ashking1):

ok so y will equal 1 right

OpenStudy (ashking1):

or am i off

OpenStudy (danjs):

8 - 16 + 9

OpenStudy (ashking1):

1

OpenStudy (danjs):

what tkhunny was doing above with completing the square, is to convert standard to vertex form, then you can read off (h,k) from that equation... That is where the -b/(2*a) can also be seen , comparing coefficients from one form to the other

OpenStudy (ashking1):

oh ok i see it now sorry many people put different method and threw me off lol